# Interlayer interactions in $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ under pressure: from $s^{\pm}$ to $d_{xy}$-wave superconductivity

### v1

@@ Abstract
For example, there has been an intense debate as to which low-energy minimal model to describe the superconducting phase. Density Functional Theory calculations [6, 7, 8, 9] reveal that multiple bands, mainly formed by Ni-centered in-plane $d_{x^{2}-y^{2}}$ and out-of-plane $d_{3z^{2}-r^{2}}$ orbitals, along with oxygen-centered $p$-orbitals cross the Fermi level, forming two electron pockets ($\alpha$, $\beta$) with mixed orbital character and one hole pocket ($\gamma$) primarily associated with the Ni-$d_{3z^{2}-r^{2}}$ orbital [6]. These bands can be described by a two-layer, two-orbital minimal tight-binding model [6, 7, 9], which has been investigated with interlayer interactions within a localized electron picture [10] and without interlayer interactions in an itinerant electron picture [11, 12, 8, 9, 13]. In all cases, RPA calculations suggest that $s^{\pm}$-wave pairing is the dominant superconducting instability, driven by intra-orbital ($\pi$, 0) scattering between the $\beta$ and the $\gamma$-pockets [9, 8]. However, the d-wave pairing channel is found to have energy close to the s-wave pairing channel [12, 8].
Other works [14, 15] propose a minimal model containing just one of these orbitals (either $d_{x^{2}-y^{2}}$ or $d_{3z^{2}-r^{2}}$), which predict an $s^{\pm}$-wave superconducting gap symmetry, with or without interlayer interactions. In contrast, a proposal of cuprate-like $d$-$p$ hybridized single orbital brought up the possibility of $d_{x^{2}-y^{2}}$ or $d_{x^{2}-y^{2}}+is$ superconductivity [16].
@@
On the experimental side, recent explorations of the ambient pressure properties of LNO327 give indications that interlayer interactions can be relevant for the superconducting phase at high pressures. Neutron scattering experiments in polycrystalline samples show evidence of spin excitations in the inelastic channel, which are consistent with strong interlayer exchange couplings [17]. In addition, X-ray absorption measurements [18] revealed dispersive magnons and spin-density wave order patterns which are also consistent with interlayer magnetic exchange being about an order of magnitude larger than intralayer ones. In this work, we investigate the role of the interlayer interaction terms in the competition between different superconducting gap symmetries in the system. We consider a two-layer, two-orbital electron model that includes both intra- and interlayer Coulomb interaction terms. Within the matrix random-phase approximation (mRPA) [19, 20, 21, 22, 23], we find that interlayer interactions favor a $d_{xy}$-wave superconductivity over the $s^{\pm}$ state, which is known to dominate when such interlayer interactions are neglected [11, 8, 14]. Our mRPA calculations reveal that interlayer interactions enhance interorbital pairing involving contributions from all three pockets, creating nodes in the gap function (absent in the $s^{\pm}$-wave state) and thus favoring the $d_{xy}$-wave pairing.
@@
Band structure.— In order to describe the (non-interacting) band structure of the system, we consider the four-band low-energy model of Refs. [6, 12]. This tight-binding model involves two orbitals ($d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$) per Ni atom in a unit cell with two Ni atoms (one for each layer) [as depicted in Fig. 1(a)] and hoppings up to second nearest neighbors. For a valence of 2.5 electrons per Ni atom [24], the resulting Fermi surface and band structure are shown in Fig. 1(b) and (c), respectively, together with the respective Ni orbital contributions.
@@
Interacting model.— Our model Hamiltonian $H=H_{0}+H_{I}$ involves both the (non-interacting) part described by $H_{0}$ as well as an interacting term ($H_{I}$) that includes both onsite ($U,J$) and interlayer ($U_{\perp},J_{\perp}$) Coulomb and exchange interaction terms, as represented in Fig. 1(a).
@@
$\begin{split}H_{\text{onsite}}&=U\sum_{ij,p\ell}n_{ip\ell\uparrow}n_{ip\ell\downarrow}\\ &+(U^{\prime}-J/2)\sum_{i\ell,p<q}n_{ip\ell}n_{iq\ell}\\ &-2J\sum_{i\ell,p<q}\bm{S}_{ip\ell}\cdot\bm{S}_{iq\ell}\\ &+J\sum_{i\ell,p<q}\left(c^{\dagger}_{ip\ell\uparrow}c^{\dagger}_{ip\ell\downarrow}c_{iq\ell\downarrow}c_{iq\ell\uparrow}+\text{h.c.}\right)\end{split}$ (1)
@@
$\begin{split}H_{\text{inter}}&=U_{\perp}\sum_{ij,p\sigma\bar{\sigma}}n_{ipA\sigma}n_{ipB\bar{\sigma}}\\ &+(U_{\perp}^{\prime}-J_{\perp}/2)\sum_{i,p<q}n_{ipA}n_{iqB}\\ &-2J_{\perp}\sum_{i,p<q}\bm{S}_{ipA}\cdot\bm{S}_{iqB}\\ &+J_{\perp}\sum_{i,p<q}\left(c^{\dagger}_{ipA\uparrow}c^{\dagger}_{ipB\downarrow}c_{iqB\downarrow}c_{iqA\uparrow}+\text{h.c.}\right).\end{split}$ (2)
In the above, $c^{\dagger}_{ip\ell\sigma}$ creates an electron with spin $\sigma$, in orbital $p$, layer $\ell=A,B$, at the (square) lattice site $i$. In addition, we use a shortcut notation for the density operator, $n_{ip\ell}=\sum_{\sigma}n_{ip\ell\sigma}$. In the present model, $U(U_{\perp})$, $U^{\prime}(U_{\perp}^{\prime})$, and $J(J_{\perp})$ are the onsite (intersite) Hubbard, interorbital, and exchange interaction strengths, respectively.
Figure 1: Schematic representation of the bilayer $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ model (a) with onsite $U,J$ and interlayer $U_{\perp},J_{\perp}$ Coulomb and exchange interaction strengths. Fermi surface (b) and band structure (c) for the two-orbital $d_{3z^{2}-r^{2}}$ (blue color) and $d_{x^{2}-y^{2}}$ (red color) low-energy model.
@@
$\begin{split}H_{I}&=\sum_{\bm{q}}\sum_{pqst\ell\ell^{\prime}}(U_{c})^{p\ell q\ell}_{s\ell^{\prime}t\ell^{\prime}}(\bm{q})n_{p\ell q\ell}(-\bm{q})n_{s\ell^{\prime}t\ell^{\prime}}(\bm{q})\\ &+\sum_{\bm{q}}\sum_{pqst\ell\ell^{\prime}}(U_{s})^{p\ell q\ell}_{s\ell^{\prime}t\ell^{\prime}}(\bm{q})\bm{S}_{p\ell q\ell}(-\bm{q})\cdot\bm{S}_{s\ell^{\prime}t\ell^{\prime}}(\bm{q}),\end{split}$ (3)
@@
$\begin{split}&(U_{s})^{pApA}_{pApA}=U,\;(U_{s})^{pApA}_{qAqA}=U^{\prime},\\ &(U_{s})^{pAqA}_{pAqA}=(U_{s})^{pAqA}_{qApA}=J,\\ &(U_{c})^{pApA}_{pApA}=U,\;(U_{c})^{pApA}_{qAqA}=(-U^{\prime}+2J),\\ &(U_{c})^{pAqA}_{qApA}=J,\;(U_{c})^{pAqA}_{pAqA}=(-J+2U^{\prime}),\\ &(U_{c})^{pApA}_{pBpB}=2U_{\perp},\;(U_{c})^{pApA}_{qBqB}=2U_{\perp}^{\prime},\\ &(U_{c})^{pAqA}_{pBqB}=(U_{c})^{pAqA}_{qBpB}=2J_{\perp},\\ \end{split}$ (4)
@@
where $A,B$ denote the respective layer index  [^1] The remaining matrix elements can be trivially obtained by exchanging $A\leftrightarrow B$. For simplicity, we assume spin rotational invariance such that the standard relationship $U^{\prime}=U-2J$ holds and we use the usual relation $J=U/4$ and $J_{\perp}=U_{\perp}/4$. Notice the interlayer interactions only play a role on the charge interaction channel.
@@
We now consider the singlet-channel RPA pairing interaction matrix [19, 21] for the multi-orbital, multilayer interacting model of Eqs. (3) and (4) [30]:
$\begin{split}\Gamma^{p\ell q\ell^{\prime}}_{s\xi t\xi^{\prime}}(\bm{q})=\frac{1}{2}&\Big{[}-\hat{U}_{c}\hat{\chi}_{c}(\bm{q})\hat{U}_{c}\\ &+3\hat{U}_{s}\hat{\chi}_{s}(\bm{q})\hat{U}_{s}+\hat{U}_{c}+\hat{U}_{s}\Big{]}^{t\xi^{\prime}q\ell^{\prime}}_{p\ell s\xi}\end{split}$ (5)
where $\hat{\chi}_{s}$ and $\hat{\chi}_{c}$ are the static spin and charge susceptibility matrices, respectively.
@@
Within the mRPA approximation, the pairing interaction $\Gamma(\bm{k},\bm{k}^{\prime})$ between Cooper pairs at $(\bm{k},-\bm{k})$ and $(\bm{k}^{\prime},-\bm{k}^{\prime})$ can be computed from Eq. (5) as [19, 20]
@@
$\begin{split}\Gamma(\bm{k},\bm{k}^{\prime})=\text{Re}&\sum_{\ell\ell^{\prime}\xi\xi^{\prime}}\sum_{pqst}\left[\psi^{t\xi^{\prime}}_{\nu_{-\bm{k}}}(-\bm{k})\right]^{*}\left[\psi^{s\xi*}_{\nu_{\bm{k}}}(\bm{k})\right]^{*}\\ &\times\Gamma^{p\ell q\ell^{\prime}}_{s\xi t\xi^{\prime}}(\bm{k}-\bm{k}^{\prime})\psi^{p\ell}_{\nu_{\bm{k}^{\prime}}}(\bm{k}^{\prime})\psi^{q\ell^{\prime}}_{\nu_{-\bm{k}^{\prime}}}(-\bm{k}^{\prime})\end{split}$ (6)
@@
where $\psi^{p\ell}_{\nu_{\bm{k}}}(\bm{k})$ is the $\nu_{\bm{k}}$-th band single-particle electronic state with momentum $\bm{k}$ in a basis where $p$ and $\ell$ are good quantum numbers. The symmetric part of the pairing interaction $(\bar{\Gamma}(\bm{k},\bm{k}^{\prime}))$ can, in turn, be used to calculate the gap function $\Delta^{\alpha}(\bm{k})$ and the respective pairing strength $\lambda^{\alpha}$ using the usual integral eigenvalue equation [19]:
@@
Figure 2: (a) Phase diagram showing the log of the ratio between the two leading eigenvalues of Eq. (7) for given values of $U$ and $U_{\perp}$. The corresponding leading gap functions $\Delta_{\rm max}(\bm{k})$ displaying $s^{\pm}$-wave and $d_{xy}$-wave symmetries for $U=1.16$ eV and $U_{\perp}/U=0.1$ (triangle) and $U_{\perp}/U=0.3$ (square) are shown in panels (b) and (c), respectively.

### v3

@@ I Introduction
For example, there has been an intense debate as to which low-energy minimal model to describe the superconducting phase. Density Functional Theory calculations [6, 7, 8, 9] reveal that multiple bands, mainly formed by Ni-centered in-plane $d_{x^{2}-y^{2}}$ and out-of-plane $d_{3z^{2}-r^{2}}$ orbitals, along with oxygen-centered $p$ orbitals cross the Fermi level, forming two electron pockets ($\alpha$, $\beta$) with mixed orbital character and one hole pocket ($\gamma$) primarily associated with the Ni $d_{3z^{2}-r^{2}}$ orbital [6]. These bands can be described by a two-layer, two-orbital minimal tight-binding model [6, 7, 9], which has been investigated with interlayer interactions within a localized electron picture [10] and without interlayer interactions in an itinerant electron picture [11, 12, 8, 9, 13].
Using these models, some RPA calculations suggest that $s^{\pm}$-wave pairing is the dominant superconducting instability [11, 14, 15, 11, 12, 16, 8], which can be driven by intra-orbital ($\pi$, 0) scattering between the $\beta$ and the $\gamma$ pockets [8], although the $d$-wave pairing channel is found to be close in energy to the $s$-wave pairing one [12, 8].
By contrast, several works have brought up the possibility that $d$-wave symmetry is dominant in some regimes. These range from $d$-$p$ hybridized single-orbital models showing cuprate-like $d_{x^{2}-y^{2}}$ symmetry [17] to two-orbital models based on Ni-$e_{g}$ DFT bands [9] as well as the differences in the spin response in these cases [18]. Moreover, a recent paper [19] reports DFT+RPA calculations showing that the leading superconducting instability has $d_{xy}$ symmetry, although a slight increase in the strength of the Ni-$e_{g}$ crystal field splitting can change the pairing symmetry to $s^{\pm}$.
@@
On the experimental side, recent explorations of the ambient pressure properties of LNO327 give indications that interlayer interactions can be relevant for the superconducting phase at high pressures. Neutron scattering experiments in polycrystalline samples show evidence of spin excitations in the inelastic channel, which are consistent with strong interlayer exchange couplings [20] while X-ray absorption measurements [21] revealed dispersive magnons and spin-density wave order patterns which are also consistent with interlayer magnetic exchange being about an order of magnitude larger than intralayer ones. In addition, recent RIXS experiments in LNO327 thin films samples show evidence of an strain-enhanced interlayer exchange interaction coupling [22].
In this work, we investigate the role of the interlayer Coulomb interaction terms in the competition between different superconducting gap symmetries in LNO327 under high pressure. We consider a two-layer, two-orbital electron model that includes both intra- and interlayer Coulomb interaction terms. Within the weak- to intermediate-coupling matrix random-phase approximation (mRPA) [23, 24, 25, 26, 27], we find that interlayer interactions favor a $d_{xy}$-wave superconductivity over the $s^{\pm}$ state, which is known to dominate when such interlayer interactions are neglected [11, 8, 14]. Our mRPA calculations reveal that interlayer interactions enhance interorbital pairing involving contributions from all three pockets, creating nodes in the gap function (absent in the $s^{\pm}$-wave state) and thus favoring the $d_{xy}$-wave pairing.
This paper is organized as follows: in Section II, we present the interacting model for bilyer LNO327, including the band structure and both the onsite and interlayer interaction terms. We review the matrix-RPA methodology and describe the paring gap calculation in Section III. Our results detailing the $s^{\pm}$-wave to $d_{xy}$-wave symmetry transition as the interlayer interaction increases are presented in Sec. IV, along with a discussion on the role of the charge susceptibility and the orbital contributions. Finally, our concluding remarks are presented in Section V.
II Model
II.1 Band structure
Figure 1: Schematic representation of the bilayer $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ model (a) with onsite $U,J$ and interlayer $U_{\perp},J_{\perp}$ Coulomb interaction and Hund’s exchange coupling strengths. FS (b) and band structure (c) for the two-orbital $d_{3z^{2}-r^{2}}$ (blue color) and $d_{x^{2}-y^{2}}$ (red color) low-energy model.
@@ II.1 Band structure
In order to describe the (non-interacting) band structure of the system, we consider the four-band low-energy model of Refs. [6, 12]. This tight-binding model involves two orbitals ($d_{x^{2}-y^{2}}$ and $d_{3z^{2}-r^{2}}$) per Ni atom in a unit cell with two Ni atoms (one for each layer) [as depicted in Fig. 1(a)] and hoppings up to second nearest neighbors. The non-interacting Hamiltonian reads
$\begin{split}H_{0}&=\sum_{\bm{k}\sigma p\ell}(T_{\bm{k}}^{p}-\mu)c_{\bm{k}p\ell\sigma}^{\dagger}c_{\bm{k}p\ell\sigma}\\ &+\sum_{\bm{k}\sigma p}t_{\perp}^{p}(c_{\bm{k}pA\sigma}^{\dagger}c_{\bm{k}pB\sigma}+\text{h.c.})\\ &+\sum_{\bm{k}\sigma(p\neq s)\ell}V_{\bm{k}}c_{\bm{k}p\ell\sigma}^{\dagger}c_{\bm{k}s\ell\sigma}\\ &+\sum_{\bm{k}\sigma(p\neq s)}V_{\bm{k}}^{\prime}(c_{\bm{k}pA\sigma}^{\dagger}c_{\bm{k}sB\sigma}+\text{h.c.}),\end{split}$ (1)
where $c^{\dagger}_{\bm{k}p\ell\sigma}$ creates an electron with spin $\sigma$, in the orbital $p=\{x,z\}$ corresponding to $\{d_{x^{2}-y^{2}},d_{3z^{2}-r^{2}}\}$, sublattice =$\ell=\{A,B\}$, and momentum $\bm{k}$, $\mu$ is the chemical potential, and h.c. shortcuts Hermitian conjugate. In the following, we use the model parameters of Ref. [6], which includes on-site energies ($\epsilon^{p}$) and up to second-neighbor hoppings such that
$\begin{split}T_{\bm{k}}^{p}&=2t_{1}^{p}(\cos k_{x}+\cos k_{y})+4t_{2}^{p}\cos k_{x}\cos k_{y}+\epsilon^{p}\;,\\ V_{\bm{k}}&=2t_{3}^{xz}(\cos k_{x}-\cos k_{y})\;,\\ V_{\bm{k}}^{\prime}&=2t_{4}^{xz}(\cos k_{x}-\cos k_{y})\;.\end{split}$ (2)
For a valence of 2.5 electrons per Ni atom [28], the resulting Fermi surface (FS) and band structure are shown in Fig. 1(b) and (c), respectively, together with the respective Ni orbital contributions.
@@
II.2 Onsite and interlayer interactions
Our model Hamiltonian $H=H_{0}+H_{I}$ involves both the non-interacting part described by $H_{0}$ as well as an interacting term ($H_{I}$) that includes both onsite ($U,J$) and interlayer ($U_{\perp},J_{\perp}$) Coulomb and Hund’s coupling interaction terms, as represented in Fig. 1(a). We remark that $J,J_{\perp}$ are the Hund’s couplings and differ from the effective exchange interaction $\tilde{J},\tilde{J}_{\perp}$ of strong-coupling models [14, 15].
@@ II.2 Onsite and interlayer interactions
$\begin{split}H_{\text{onsite}}&=U\sum_{ij,p\ell}n_{ip\ell\uparrow}n_{ip\ell\downarrow}\\ &+(U^{\prime}-J/2)\sum_{i\ell,p<q}n_{ip\ell}n_{iq\ell}\\ &-2J\sum_{i\ell,p<q}\bm{S}_{ip\ell}\cdot\bm{S}_{iq\ell}\\ &+J^{\prime}\sum_{i\ell,p<q}\left(c^{\dagger}_{ip\ell\uparrow}c^{\dagger}_{ip\ell\downarrow}c_{iq\ell\downarrow}c_{iq\ell\uparrow}+\text{h.c.}\right)\end{split}$ (3)
@@
$\begin{split}H_{\text{inter}}&=U_{\perp}\sum_{ij,p\sigma\bar{\sigma}}n_{ipA\sigma}n_{ipB\bar{\sigma}}\\ &+(U_{\perp}^{\prime}-J_{\perp}/2)\sum_{i,p<q}n_{ipA}n_{iqB}\\ &-2J_{\perp}\sum_{i,p<q}\bm{S}_{ipA}\cdot\bm{S}_{iqB}\\ &+J^{\prime}_{\perp}\sum_{i,p<q}\left(c^{\dagger}_{ipA\uparrow}c^{\dagger}_{ipB\downarrow}c_{iqB\downarrow}c_{iqA\uparrow}+\text{h.c.}\right).\end{split}$ (4)
In the above, $c^{\dagger}_{ip\ell\sigma}$ creates an electron with spin $\sigma$, in orbital $p$, layer $\ell=A,B$, at the (square) lattice site $i$. In addition, we use a shortcut notation for the density operator, $n_{ip\ell}=\sum_{\sigma}n_{ip\ell\sigma}$. In the present model, $U(U_{\perp})$, $U^{\prime}(U_{\perp}^{\prime})$, $J(J_{\perp})$, and $J^{\prime}(J^{\prime}_{\perp})$ are the onsite (intersite) Hubbard, interorbital Hubbard, Hund’s exchange coupling and pair-hopping interaction strengths, respectively. In the following, we assume local spin rotational invariance such that $J^{\prime}\!=\!J(J^{\prime}_{\perp}\!=\!J_{\perp})$ [23, 24].
@@
$\begin{split}H_{I}&=\sum_{\bm{q}}\sum_{pqst\ell\ell^{\prime}}(U_{c})^{p\ell q\ell}_{s\ell^{\prime}t\ell^{\prime}}n_{p\ell q\ell}(-\bm{q})n_{s\ell^{\prime}t\ell^{\prime}}(\bm{q})\\ &+\sum_{\bm{q}}\sum_{pqst\ell\ell^{\prime}}(U_{s})^{p\ell q\ell}_{s\ell^{\prime}t\ell^{\prime}}\bm{S}_{p\ell q\ell}(-\bm{q})\cdot\bm{S}_{s\ell^{\prime}t\ell^{\prime}}(\bm{q}),\end{split}$ (5)
@@
$\begin{split}&(U_{s})^{pApA}_{pApA}=U,\;(U_{s})^{pApA}_{qAqA}=U^{\prime},\\ &(U_{s})^{pAqA}_{pAqA}=(U_{s})^{pAqA}_{qApA}=J,\\ &(U_{c})^{pApA}_{pApA}=U,\;(U_{c})^{pApA}_{qAqA}=-U^{\prime}+2J,\\ &(U_{c})^{pAqA}_{qApA}=J,\;(U_{c})^{pAqA}_{pAqA}=-J+2U^{\prime},\\ &(U_{c})^{pApA}_{pBpB}=2U_{\perp},\;(U_{c})^{pApA}_{qBqB}=2U^{\prime}_{\perp},\\ &(U_{c})^{pAqA}_{pBqB}=(U_{c})^{pAqA}_{qBpB}=2J_{\perp},\\ \end{split}$ (6)
@@
where $A,B$ denote the respective layer index  [^1] The remaining matrix elements can be trivially obtained by exchanging $A\leftrightarrow B$. For simplicity, we assume spin rotational invariance such that the standard relationships $U^{\prime}=U-2J$ and $U^{\prime}_{\perp}=U_{\perp}-2J_{\perp}$ hold and we set $J=U/4$ and $J_{\perp}=U_{\perp}/4$, which is consistent with previous studies [7], although we have verified that this choice does not qualitatively alter our main results (see Appendix A). Notice that the interlayer interaction terms $(U_{\perp},U^{\prime}_{\perp})$ appear only in the charge interaction matrices.
III Matrix RPA method
III.1 Spin and charge susceptibilities
For completeness, we begin this Section by briefly reviewing the RPA-enhanced spin and charge susceptibilities for the multiorbital model of Eq. (1). Following Ref. [23], the bare spin susceptibility is calculated from the spin-spin correlation function as
[…]
where $i\nu_{m}$ denotes Matsubara bosonic frequencies, $\tau$ is the imaginary time, $T$ is the temperature, and $T_{\tau}$ is the time-ordering operator. For the non-interacting model, the spin and charge susceptibilities are equivalent [23] and can be written in terms of the non-interacting Green’s functions as
$\begin{split}[\hat{\chi}]^{p\ell q\ell^{\prime}}_{s\xi t\xi^{\prime}}(\bm{q},i\nu_{m})=&-\frac{T}{N}\sum_{\bm{k}i\omega_{n}}G_{s\xi p\ell}(\bm{k},i\omega_{n})\\ &\times G_{q\ell^{\prime}t\xi^{\prime}}(\bm{k}+\bm{q},i\omega_{n}+i\nu_{m}),\end{split}$ (8)
where
$G_{s\xi p\ell}(\bm{k},i\omega_{n})=\sum_{\nu_{\bm{k}}}\frac{\psi^{s\xi}_{\nu_{\bm{k}}}(\bm{k})\psi^{p\ell*}_{\nu_{\bm{k}}}(\bm{k})}{i\omega_{n}-E_{\nu_{\bm{k}}}(\bm{k})},$ (9)
and $\psi^{p\ell}_{\nu_{\bm{k}}}(\bm{k})$ is a single-particle eigenfunction of $H_{0}$ (given by Eq. (1)) related to the $\nu_{\bm{k}}$-th electronic band with energy $E_{\nu_{\bm{k}}}(\bm{k})$, and written in a basis where $p$ and $\ell$ are good quantum numbers. The Matsubara summation is calculated numerically by using Ozaki’s method [34]. In this work, we adopt $T=0.02$ eV.
$\hat{\chi}_{s}(\bm{q},i\nu_{m})$ $=\hat{\chi}(\bm{q},i\nu_{m})[\hat{1}-\hat{U}_{s}\hat{\chi}(\bm{q},i\nu_{m})]^{-1},$ (10)
$\hat{\chi}_{c}(\bm{q},i\nu_{m})$ $=\hat{\chi}(\bm{q},i\nu_{m})[\hat{1}+\hat{U}_{c}\hat{\chi}(\bm{q},i\nu_{m})]^{-1}.$ (11)
where $\hat{\chi}(\bm{q},i\nu_{m})$ is the bare susceptibility matrix given by Eq. (8). We have checked that the Stoner criterion is not matched for the range of interaction strengths studied in the present work.
Figure 2: Contour plots for the main eigenvalue of the spin (a) and charge (b,c,d) RPA susceptibility matrices for $U=1.16$ eV and different $U_{\perp}/U$ ratios.
It is important to notice that $\hat{U}_{s}$ does not depend on $U_{\perp}$ while $\hat{U}_{c}$ does (see Eq. (6)). As such, only the charge susceptibility changes as the interlayer interactions increase. This is illustrated in Figure 2, which shows the main eigenvalue of the spin [$\hat{\chi}_{s}(\bm{q},0)$] and charge [$\hat{\chi}_{c}(\bm{q},0)$] static susceptibility matrices for different values of $U_{\perp}/U$ for $U=1.16$ eV. As such, $\hat{\chi}_{s}(\bm{q})$ does not change as $U_{\perp}/U$ varies, while the main changes in $\hat{\chi}_{c}(\bm{q})$ are a reduction near the $(\pi,\pi)$ ($\bm{M}$) point and and overall reduction around the $(0,0)$ ($\bm{\Gamma}$) point. These features will be further discussed in Section IV.2.
III.2 RPA pairing interaction
@@
We now consider the singlet-channel RPA pairing interaction matrix [23, 25] for the multi-orbital, multilayer interacting model of Eqs. (5) and (6)
$\begin{split}\Gamma^{p\ell q\ell^{\prime}}_{s\xi t\xi^{\prime}}(\bm{q})=\frac{1}{2}\Big{[}&-\hat{U}_{c}\hat{\chi}_{c}(\bm{q})\hat{U}_{c}\\ &+3\hat{U}_{s}\hat{\chi}_{s}(\bm{q})\hat{U}_{s}+\hat{U}_{c}+\hat{U}_{s}\Big{]}^{t\xi^{\prime}q\ell^{\prime}}_{p\ell s\xi}.\end{split}$ (12)
@@
In the mRPA, the pairing interaction $\Gamma(\bm{k},\bm{k}^{\prime})$ between Cooper pairs at $(\bm{k},-\bm{k})$ and $(\bm{k}^{\prime},-\bm{k}^{\prime})$ is computed from Eq. (12) as [23, 24]
@@
$\begin{split}\Gamma(\bm{k},\bm{k}^{\prime})=\text{Re}&\sum_{\ell\ell^{\prime}\xi\xi^{\prime}}\sum_{pqst}\left[\psi^{t\xi^{\prime}}_{\nu_{-\bm{k}}}(-\bm{k})\right]^{*}\left[\psi^{s\xi*}_{\nu_{\bm{k}}}(\bm{k})\right]^{*}\\ &\times\Gamma^{p\ell q\ell^{\prime}}_{s\xi t\xi^{\prime}}(\bm{k}-\bm{k}^{\prime})\psi^{p\ell}_{\nu_{\bm{k}^{\prime}}}(\bm{k}^{\prime})\psi^{q\ell^{\prime}}_{\nu_{-\bm{k}^{\prime}}}(-\bm{k}^{\prime})\;.\end{split}$ (13)
@@
The symmetric part of the pairing interaction $2\bar{\Gamma}(\bm{k},\bm{k}^{\prime})={\Gamma}(\bm{k},\bm{k}^{\prime})+{\Gamma}(\bm{k},-\bm{k}^{\prime})$ is, in turn, used to calculate the gap function $\Delta^{\alpha}(\bm{k})$ and the respective pairing strength $\lambda^{\alpha}$ using the usual integral eigenvalue equation [23]
@@
Figure 3: (a) Phase diagram showing the log of the ratio between the two leading eigenvalues of Eq. (14) for given values of $U$ and $U_{\perp}$. The corresponding leading gap functions $\Delta_{\rm max}(\bm{k})$ displaying $s^{\pm}$-wave and $d_{xy}$-wave symmetries for $U=1.16$ eV and $U_{\perp}/U=0.1$ ($\triangle$) and $U_{\perp}/U=0.3$ ($\square$) are shown in panels (b) and (c), respectively.
IV Superconducting paring symmetries
IV.1 Leading solutions of the gap equation
For a given set of parameters entering $\bar{\Gamma}(\bm{k},\bm{k}^{\prime})$, the solution of Eq. (14) with the highest eigenvalue $\lambda_{\rm max}$ will correspond to the highest critical temperature $T_{c}$. The symmetry of the pairing gap is reflected in the corresponding eigenfunction $\Delta_{\rm max}(\bm{k})$.